A hockey puck (1) of mass 120 g is shot east at a speed of 8.80 m/s. It strikes a second puck (2), initially at rest, of mass 138 g. As a result of the collision, the first puck (1) is deflected at an angle of 46° south of east and the second puck (2) moves at an angle of 40° north of east. What is the magnitude of the velocity of puck (1) after the collision?
Answer #1

yunus mu,+ m2 = mivienot my coma (012) (88-8) to: 138(0) = (0-120) V. But +0138/2 2014 -0 1.056= 0.0833v, + 0.1057 v2 alons y